Petal:

The map of a walking trail is drawn on a coordinate grid with three points of interest. The trail starts at R(−2, 4) and goes to S(3, 4) and continues to T(3, −1). The total length of the walking trail is?

jacob28:

R(-2, 4) to S(3, 4) is sqrt((3 - (-2))^2 + (4 - 4)^2) = sqrt(5)^2 = 5 units S(3, 4) to T(3, -1) is sqrt((3 - 3)^2 + (-1 - 4)^2) = sqrt(-5)^2 = 5 units So what will the total length be?

jacob28:

ye thankks